OpenStudy (anonymous):

It takes Brian 15 hours longer to build a model car than it takes John. If they work together, they can build the model car in 4 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Brian to build the car on his own.

OpenStudy (anonymous):

someone please help

OpenStudy (boldjon):

1/x + 1/(x + 15) = 1/4 1 car per x hours + 1 car per (x+15) hours equals 1 car in 4 hours. Solve for x. x = 5 Brian takes 5 + 15 = 20 hours

OpenStudy (anonymous):

you would have to know how long it takes for john to build a model car and then you do x + 15 + x = y. If x = 10 then the answer is 35

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

can someone show me

OpenStudy (anonymous):

nvm it would be 25

OpenStudy (anonymous):

I've tried so many times and failed. i know brian can complete it in 20 hours

OpenStudy (anonymous):

I messed up in calculation but you would add x to 15 x = (how long john takes to build a model car.) if it took brian 20 hours it would mean it took john 5 hours you see what I mean you add x to 15 and that is the answer

OpenStudy (anonymous):